OpenStudy (anonymous):

Say you have a right triangle with side a = 40 and side c = 58. What is the length of side b

OpenStudy (anonymous):

40^2 + x^2 = 58^2 x=42 ---> length of side b

OpenStudy (anonymous):

So here you use pythagorean theorem a^2 + b^2 = c^2 you are given a= 40 and c= 58 b^2 = c^2 - a^2 plug the numbers in and you will get your answer :)

OpenStudy (anonymous):

does ^ mean squared?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[a^{2} + b ^{2} = c ^{2} \]

OpenStudy (anonymous):

but i dont know what B is

OpenStudy (anonymous):

but you know a and c so there is only 1 unknown variable so rearrange the formula to b^2 = c^2 - a^2

OpenStudy (anonymous):

\[b= \sqrt{c ^{2}-a^{2}}\]

OpenStudy (anonymous):

to square a number u times it by itself or times 2?

OpenStudy (anonymous):

times it by itself

OpenStudy (anonymous):

3^2 = 9 < 3x3 but 3x2 = 6